HUBE 466-566-P12020(2) - Compatibility Mode 18109361685 0 - 0 X File Home Insert Draw Design Layout References Mailings RevieI need help with problem 3 only, please show all the calculations with details, don't just list numbers or put them on a table using excel. I need to understand how did you get each thing. Thanks
Answer #1
In the given problem Home Currency is USD($).
Therefore,
For Direct Quote
Relationship Appreciation Depreciation
Forward Rate(F) > Spot Rate(S) Foreign Currency Home Currency
Forward Rate(F) < Spot Rate(S) Home Currency Foreign Currency
For Inirect Quote
Relationship Appreciation Depreciation
Forward Rate(F) > Spot Rate(S) Home Currency Foreign Currency
Forward Rate(F) < Spot Rate(S) Foreign Currency Home Currency
Appreciation means Currency is at Premium and Depreciation means
Currency is at Discount.
Formula for Premium or Discount
In case of Direct Quote = (F-S)/S x 12/n x 100
In case of Indirect Quote = (S-F)/F x 12/n x 100
Where,
F = Forward Rate
S = Spot Rate
n = No of Months of Forward Conntract.
a. Euro
Spot Rate (S) = $1.1052 / €
Forward Rate (F) = $1.2560 / €
Home Currency is USD so this is a Direct Quote.
As per Direct Quote, when F>S Foreign Currency i.e. Euro is appreciating and Home
Currency i.e. USD is depreciating.
That means Euro is at premium and USD is at discount in 1 year Forward Market.
As per Direct Quote,
% of Forward Discount = (F-S)/S x 12/n x 100
=($1.2560 - $1.1052)/$1.1052 x 12/12 x 100
=13.64%
b. Mexican Peso (MXN)
Spot Rate (S) = MXN19.5905/$
Forward Rate (F) = MXN22.2775/$
Home Currency is USD so this is a Indirect Quote.
As per Indirect Quote, when F>S Home Currency i.e. USD is appreciating and Foreign
Currency i.e. Mexican Peso is depreciating.
That means USD is at premium and Mexican Peso is at discount in 1 year Forward Market.
As per Indirect Quote,
% of Forward Premium = (S-F)/F x 12/n x 100
=(MXN19.5905 - MXN22.2775)/MXN22.2775 x 12/12 x 100
=12.06%
c. British Pound
Spot Rate (S) = $1.3350/£
Forward Rate (F) = $1.3020/£
Home Currency is USD so this is a Direct Quote.
As per Direct Quote, when F<S Home Currency i.e. USD is appreciating and Foreiegn
Currency i.e. British Pound is depreciating.
That means USD is at premium and British Pound is at discount in 1 year Forward Market.
As per Direct Quote,
% of Forward Premium = (F-S)/S x 12/n x 100
=($1.3020 - $1.3350)/$1.3350 x 12/12 x 100
=2.47%
d. Japanese Yen
Spot Rate (S) = ¥109.2802/$
Forward Rate (F) = ¥116.3225/$
Home Currency is USD so this is a Indirect Quote.
As per Indirect Quote, when F>S Home Currency i.e. USD is appreciating and Foreign
Currency i.e. Japanese Yen is depreciating.
That means USD is at premium and Japanese Yen is at discount in 1 year Forward Market.
As per Indirect Quote,
% of Forward Premium = (S-F)/F x 12/n x 100
=(¥109.2802 - ¥116.3225)/¥116.3225 x 12/12 x 100
=6.05%